# CryoFM: A Flow-based Foundation Model for Cryo-EM Densities

### v1

@@ 4.1 Pretraining Dataset
Our training dataset consists of deposited sharpened density maps from the EMDB (wwPDB Consortium, 2023), specifically those with: 1) a reported resolution better than $3.0$, 2) structures resolved by single-particle cryo-EM, and 3) data entries that include half-maps, ensuring that resolution estimates are based on the “gold standard” Fourier shell correlation (FSC) (Henderson et al., 2012). We manually curated this subset by removing exceptionally large complexes, helical structures and problematic cases through visual inspection. We also excluded density maps with side lengths greater than $576$. This curation resulted in a total of $3479$ density maps for model training and testing. The density maps in the selected subset were lowpass filtered to $1.5$/voxel and $3$/voxel for CryoFM-S and CryoFM-L, respectively. A total of $32$ density maps were selected as test set and excluded from training. For CryoFM-S, we applied random cropping to volumes of size $64^{3}$ along with random rotations for augmentation, whereas data for CryoFM-L were center cropped to volumes of size $128^{3}$ and augmented solely with random rotations.
@@ 4.2 Architecture
A major challenge in applying the vanilla Transforemr architecture to 3D density ${\mathbf{x}}_{t}\in\mathbb{R}^{D^{3}}$ is the computational complexity, which is $\mathcal{O}(D^{6})$  [^5] The computational complexity of Transformers is $\mathcal{O}(L^{2})$, where $L=D^{3}$ is the number of tokens. . CryoFM processes the 3D input data through a Transformer with a hierarchical architecture, based on HDiT (Crowson et al., 2024). As illustrated in Fig. 2, the hierarchical structure downsamples the spatial dimension at initial levels and upsamples it at final levels, implemented via PixelUnshuffle and PixelShuffle layers (Shi et al., 2016). Consequently, the spatial dimension at the middle level is significantly reduced, making the model more easily scalable. Furthermore, at the two ends of the hierarchical structure, CryoFM employs Neighborhood Attention (NA) (Hassani et al., 2023) since a localized attention layer is able to capture local dependencies while considerably reducing the number of tokens to attend.
Additionally, it is important to note that we normalize the data ${\mathbf{x}}_{0}$ before training CryoFM to avoid large changes in the variance of model’s output. We empirically sample some data from the dataset and calculate a mean value of $0.04$ and a standard deviation of $0.09$.
@@ 4.3 Flow Posterior Sampling
Given a vector field $v_{t}({\mathbf{x}}_{t})$ that generates the prior distribution $p_{0}({\mathbf{x}}_{0})$, we aim to convert it to a vector field $v_{t}({\mathbf{x}}_{t}|{\mathbf{y}})$ that generates the posterior $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$. We firstly correlate the vector field in Eq. (1) with the score function $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ in that (Dao et al., 2023; Song et al., 2021b):
$v_{t}({\mathbf{x}}_{t})=f_{t}({\mathbf{x}}_{t})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}),$
where $f_{t}$ is the drift term and $g_{t}$ is the diffusion coefficient in the forward-time SDE [^6] An SDE of the standard form is $d{\mathbf{x}}_{t}=f_{t}({\mathbf{x}}_{t})dt+g_{t}d{\mathbf{w}}$, where $d{\mathbf{w}}$ is the Wiener process. The probability ODE which has the same marginals as the reverse of the SDE is $d{\mathbf{x}}_{t}=\left(f_{t}({\mathbf{x}}_{t})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\right)dt$. Compared with Eq. (1), we can find the relationship between the vector field and the score function.. Plugging the condition into both sides, we have:
$v_{t}({\mathbf{x}}_{t}|{\mathbf{y}})=f_{t}({\mathbf{x}}_{t}|{\mathbf{y}})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{y}})=v_{t}({\mathbf{x}}_{t})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t}).$
Thus, adding a likelihood term weighted by the diffusion coefficient generates a vector field that models the posterior distribution. We follow Dao et al. (2023) to set $f_{t}({\mathbf{x}}_{t})=-\frac{{\mathbf{x}}_{t}}{1-t}$ and $\frac{g_{t}^{2}}{2}=\frac{t}{1-t}$ for the flow trajectory defined in Eq. (2), so the conditional vector field is:
$v_{t}({\mathbf{x}}_{t}|{\mathbf{y}})=v_{t}({\mathbf{x}}_{t})-\frac{t}{1-t}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t}).$
Similar to Section 3.2, we estimate the likelihood term by:
$v_{t}({\mathbf{x}}_{t}|{\mathbf{y}})\approx v_{t}({\mathbf{x}}_{t})+\frac{t}{1-t}\cdot\lambda_{t}\nabla_{{\mathbf{x}}_{t}}\|{\mathbf{y}}-\mathcal{A}\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}_{2}.$
In practice, we incorporate some tricks to avoid numerical instability. The detailed algorithm of flow posterior sampling is explained in Alg. 1. The algorithm can be adapted for different tasks; we refer the reader to Appendix E for the additional versions.
@@
a pretrained vector field $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, a degradation operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, an obseravation ${\mathbf{y}}\in\mathbb{R}^{m}$, number of steps $N$, maximum step size $\lambda_{\text{max}}$ at each step
@@
for $t\in[1,1-\frac{1}{N},1-\frac{2}{N},\cdots,0]$ do
     ${\mathbf{x}}^{\prime}_{t-1}\leftarrow{\mathbf{x}}_{t}-v_{\theta}(t,{\mathbf{x}}_{t})$
@@
     ${\mathbf{x}}_{t-1}\leftarrow{\mathbf{x}}^{\prime}_{t-1}+\lambda_{t}{\mathbf{g}}$
@@ 5 Experiments
Fourier Shell Correlation (FSC) is a widely used metric that compares two density maps in Fourier space (Harauz & van Heel, 1986), allowing for the assessment of alignment between the ground truth and the reconstructed map. In the experiments, to evaluate the quality of the density maps, we use three primary metrics: $\text{FSC}_{\text{AUC}}$, $\text{FSC}_{0.5}$, and Fail Rate (FR). Specifically, $\text{FSC}_{\text{AUC}}$ measures the overall correlation across all spatial frequencies, and $\text{FSC}_{0.5}$ represents the resolution of the reconstructed map at the standard 0.5 cutoff (Rosenthal & Henderson, 2003). The Fail Rate (FR) identifies cases where the method fails to run or produces a result that significantly deviates from the ground truth. For the FSC metrics, we only report the results for cases that did not fail. More detailed explanation can be found in Section C.1.
@@ 5.1 Spectral Noise Denoising
In cryo-EM reconstruction, spectral noise is the most commonly used noise model due to its ability to capture the varying noise characteristics across different spatial frequencies, which arise from factors such as the contrast transfer function (CTF) and detector imperfections. We consider the introduction of noise in the Fourier domain, where the variance of the added Gaussian noise is frequency-dependent, with higher frequencies exhibiting greater noise variance. Given a density $\tilde{{\mathbf{V}}}\in\mathbb{C}^{D\times D\times D}$ in the Fourier domain, the degradation model $\mathcal{A}:\mathbb{C}^{D\times D\times D}\to\mathbb{C}^{D\times D\times D}$ is:
@@
In the experiment, we estimate $\sigma_{\text{noise}}^{2}(\nu)$ from the half maps, which are two independent observations with the same underlying signal with uncorrelated noise. See Section D.1 for details.
Table 1: Results of the spectral noise denoising task, comparing CryoFM with DeepEMhancer (Sanchez-Garcia et al., 2021) and EMReady (He et al., 2023). The best attainable $\text{FSC}_{0.5}$ is $3.0$ since the voxel size is $1.5$.
@@
Figure 3: Result of the spectral noise denoising task. Two density maps from EMDB were added spectral noise so that the estimated resolution is $4.3$. The degraded density maps were filtered by relion_postprocess for visual clarity.
We manually introduce spectral noise, resulting in resolutions ranging from 3.2 Å to 15.0 Å. DeepEMhancer and EMReady, both pretrained models that directly learn the posterior $p({\mathbf{x}}|{\mathbf{y}})$, are used as baselines. CryoFM demonstrates superior robustness, with no failed cases across all experiments, while both baselines show higher fail rates, especially at medium resolutions. Additionally, as shown in Tab. 1 and Tab. 5 in Section D.1, CryoFM achieves higher $\text{FSC}_{\text{AUC}}$ and better $\text{FSC}_{0.5}$ in most cases, significantly improving degraded density maps and enhancing the resolution of helices and loops, as illustrated in Fig. 3. Notably, the FSC curves demonstrate that CryoFM successfully improves the signal across all frequencies, while the baselines fails to retain trustworthy signals in the low-frequency range.
@@ 5.2 Anisotropic Noise Denoising
Figure 4: Degradation operators for (i) the anisotropic noise and (ii) the missing wedge.
Anisotropic noise in cryo-EM occurs when noise distribution varies by direction, affecting some orientations more than others and leading to uneven reconstruction quality. As a result, certain orientations have lower signal-to-noise ratios than others. To approximate this degradation in a simplified manner, we amplify the spectral noise by a factor when the particle orientation falls within a specific range of angles, simulating the increased uncertainty in less-sampled orientations. The degradation model $\mathcal{A}:\mathbb{C}^{D\times D\times D}\to\mathbb{C}^{D\times D\times D}$ is given by:
@@
where ${\mathbf{\epsilon}}\in\mathbb{R}^{D\times D\times D}$ is the same as defined in Eq. (3), and $\alpha>1$ is a scaler that increases the noise for specific orientation angles, reflecting the heightened uncertainty in those directions. $\theta_{\min}$ and $\theta_{\max}$ are the angles that controls the portion of the noise being amplified as illustrated in Fig. 4. In the experiment, we estimate $\sigma_{\text{noise}}^{2}(\nu)$ and $\alpha$ from the half maps. See Section D.2 for details.
@@
Figure 5: Result of the anisotropic noise denoising task. Two density maps from EMDB were added anisotropic spectral noise that the estimated resolution is 4.38 Å. The degraded density maps were filtered by relion_postprocess for visual clarity.
@@ 5.3 Missing Wedge Restoration
The missing wedge problem in cryo-ET arises from the limited tilt range of the electron microscope during data acquisition, causing incomplete Fourier space sampling and artifacts in the subtomograms. We simulate this effect by applying a wedge-shaped mask in the Fourier domain, removing data from unmeasured orientations. The degradation model $\mathcal{A}:\mathbb{C}^{D\times D\times D}\to\mathbb{C}^{D\times D\times D}$ is given by:
@@
where $\theta_{\min}$ and $\theta_{\max}$ are typically set to $-60^{\circ}$ and $+60^{\circ}$ respectively, representing the common tilt angle limits in experimental setups as illustrated in Fig. 4. For posterior sampling, we slightly modify the flow posterior sampling algorithm to Alg. 2.
@@
Figure 6: Result of missing wedge restoration task. Two density maps from EMDB were masked in Fourier space to simulate the missing wedge effect in cryo-ET. The central slices through the maps were also shown.
@@ 5.4 Ab initio Modeling
Ab initio modeling in cryo-EM involves reconstructing a coarse density map from 2D particle projections, which serves as a reference for refinement. Here, we simplify the problem by focusing on generating a coarse density map from a few clean 2D projections obtained via upstream 2D classification. This approach mirrors earlier ab initio methods (Ludtke et al., 1999; Voss et al., 2010) before cryoSPARC’s introduction of SGD (Punjani et al., 2017). Given $K$ projections, there exist $K$ degradation operators, where the $k$-th degradation model is $\mathcal{A}^{(k)}:\mathbb{R}^{D\times D\times D}\to\mathbb{R}^{D\times D}$ given by:
@@
where $\mathcal{P}$ is a projection operator in the real space, and $\phi^{(k)}\in\mathbb{SO}(3)\times\mathbb{R}^{2}$ is the pose of the 2D projections. In this task, $\phi^{(k)}$ is unknown and can not be pre-determined, so it must be iteratively searched to find the optimal value (Scheres, 2012a; Punjani et al., 2017; Zhong et al., 2021b), we modify the flow posterior sampling to Alg. 3. Moreover, since ab initio modeling focuses on capturing the global shape at low resolution, we use CryoFM-L for this task.

### v2

@@ 4.1 Pretraining Dataset
Our training dataset consists of deposited sharpened density maps from the EMDB (wwPDB Consortium, 2023a), specifically those with: 1) a reported resolution better than $3.0$, 2) structures resolved by single-particle cryo-EM, and 3) data entries that include half-maps, ensuring that resolution estimates are based on the “gold standard” Fourier shell correlation (FSC) (Henderson et al., 2012). We manually curated this subset by removing exceptionally large complexes, helical structures, and problematic cases through visual inspection  (see Fig. 13 for examples). We also excluded density maps with side lengths greater than $576$. This curation resulted in a total of $3479$ density maps, where $32$ density maps were selected as test set and excluded from training. The density maps in the selected subset were lowpass filtered to $1.5$/voxel and $3$/voxel for CryoFM-S and CryoFM-L, respectively. For CryoFM-S, we applied random cropping to volumes of size $64^{3}$ along with random rotations for augmentation, whereas data for CryoFM-L were center cropped to volumes of size $128^{3}$ and augmented solely with random rotations. Additionally, we rescale the density value to avoid large numerical variance  (Section C.2 describes the details).
@@ 4.2 Architecture
A major challenge in applying the vanilla Transformer architecture to 3D density ${\mathbf{x}}_{t}\in\mathbb{R}^{D^{3}}$ is the computational complexity, which is $\mathcal{O}(D^{6})$  [^4] The computational complexity of Transformers is $\mathcal{O}(L^{2})$, where $L=D^{3}$ is the number of tokens. . CryoFM processes the 3D input data through a Transformer with a hierarchical architecture, based on HDiT (Crowson et al., 2024). As illustrated in Fig. 2, the hierarchical structure downsamples the spatial dimension at initial levels and upsamples it at final levels, implemented via PixelUnshuffle and PixelShuffle layers (Shi et al., 2016). Consequently, the spatial dimension at the middle level is significantly reduced, making the model more easily scalable. Furthermore, at the two ends of the hierarchical structure, CryoFM employs Neighborhood Attention (NA) (Hassani et al., 2023) since a localized attention layer is able to capture local dependencies while considerably reducing the number of tokens to attend.
@@ 4.3 Flow Posterior Sampling
Given a vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$ that generates the prior data distribution $p_{0}({\mathbf{x}}_{0})$, we aim to convert it to a vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}|{\mathbf{y}})$ that generates the posterior $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$. First, according to Dao et al. (2023); Song et al. (2021b), we can connect the vector field with the score function $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ as follows:
${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})=f_{t}({\mathbf{x}}_{t})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}),$
where $f_{t}$ is the drift term and $g_{t}$ is the diffusion coefficient in the forward-time SDE. In addition, we define the conditional vector field as follows:
${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}|{\mathbf{y}})=f_{t}({\mathbf{x}}_{t})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{y}})={\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})-\frac{g_{t}^{2}}{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t}),$
where the second equality holds due to Eq. (3). Thus, substracting the score of the likelihood term weighted by the diffusion coefficient generates a vector field that models the posterior distribution. Following the rectified flow (Dao et al., 2023), we set $f_{t}({\mathbf{x}}_{t})=-\frac{{\mathbf{x}}_{t}}{1-t}$ and $\frac{g_{t}^{2}}{2}=\frac{t}{1-t}$ in the conditional flow in Eq. (2), and obtain the conditional vector field as follows:
${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}|{\mathbf{y}})={\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})-\frac{t}{1-t}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t}).$ (5)
Similar to Section 3.2, by substituting Eq. (4) into Eq. (5), we obtain:
${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}|{\mathbf{y}})\approx{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})+\frac{t}{1-t}\cdot\lambda_{t}\nabla_{{\mathbf{x}}_{t}}\|{\mathbf{y}}-\mathcal{A}\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}_{2}.$
where $\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})={\mathbf{x}}_{t}-t\cdot{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$ is the posterior mean estimated by the flow model. In practice, we incorporate some tricks to avoid numerical instability. The detailed algorithm of flow posterior sampling is described in Alg. 1. The algorithm can be adapted for different tasks; we refer the reader to Appendix G for the additional versions.
@@
a pretrained vector field ${\textnormal{v}}_{\Theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, a forward operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, an obseravation ${\mathbf{y}}\in\mathbb{R}^{m}$, number of steps $N$, maximum step size $\lambda_{\text{max}}$ at each step
@@
for $t\in[1,1-\Delta t,1-2\Delta t,\cdots,\Delta t]$ do
     ${\mathbf{x}}^{\prime}_{t-\Delta t}\leftarrow{\mathbf{x}}_{t}-\Delta t\cdot{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$
@@
     ${\mathbf{x}}_{t-\Delta t}\leftarrow{\mathbf{x}}^{\prime}_{t-\Delta t}-\lambda_{t}{\mathbf{g}}$
@@ 5 Experiments
Fourier Shell Correlation (FSC) is a widely used metric that compares two density maps in Fourier space (Harauz & van Heel, 1986), allowing for the assessment of alignment between the ground truth and the reconstructed map. In the experiments, to evaluate the quality of the density maps, we use three primary metrics: $\text{FSC}_{\text{AUC}}$, $\text{FSC}_{0.5}$, and Fail Rate (FR). Specifically, $\text{FSC}_{\text{AUC}}$ measures the overall correlation across all spatial frequencies, and $\text{FSC}_{0.5}$ represents the resolution of the reconstructed map at the standard 0.5 cutoff (Rosenthal & Henderson, 2003). The Fail Rate (FR) identifies cases where the method fails to run or produces a result that significantly deviates from the ground truth. For the FSC metrics, we only report the results for cases that did not fail. More detailed explanation can be found in Section E.1.
@@ 5.1 Spectral Noise Denoising
In cryo-EM reconstruction, spectral noise is the most commonly used noise model due to its ability to capture the varying noise characteristics across different spatial frequencies, which arise from factors such as the contrast transfer function (CTF) and detector imperfections. We consider the introduction of noise in the Fourier domain, where the variance of the added Gaussian noise is frequency-dependent, with higher frequencies exhibiting greater noise variance. Given a density $\tilde{{\mathbf{V}}}\in\mathbb{C}^{D\times D\times D}$ in the Fourier domain, the forward model $\mathcal{A}:\mathbb{C}^{D\times D\times D}\to\mathbb{C}^{D\times D\times D}$ is:
@@
In the experiment, we estimate $\sigma_{\text{noise}}^{2}(\nu)$ from the half maps, which are two independent observations with the same underlying signal with uncorrelated noise. See Section F.1 for details.
Table 1: Results of the spectral noise denoising task, comparing CryoFM with DeepEMhancer (Sanchez-Garcia et al., 2021) and EMReady (He et al., 2023). In Sections 5.1, 5.2, 5.3, the forward model acts as a degradation of the original data. The best attainable $\text{FSC}_{0.5}$ is $3.0$ since the voxel size is $1.5$.
@@
Figure 3: Result of the spectral noise denoising task. Two density maps from EMDB were added spectral noise so that the estimated resolution is $4.3$. The degraded density maps (results after applying the forward model) were filtered by relion_postprocess for visual clarity.
We manually introduce spectral noise, resulting in resolutions ranging from 3.2 Å to 15.0 Å. DeepEMhancer and EMReady, both pretrained models that directly learn the posterior $p({\mathbf{x}}|{\mathbf{y}})$, are used as baselines. CryoFM demonstrates superior robustness, with no failed cases across all experiments, while both baselines show higher fail rates, especially at medium resolutions. Additionally, as shown in Tab. 1 and Tab. 5 in Section F.1, CryoFM achieves higher $\text{FSC}_{\text{AUC}}$ and better $\text{FSC}_{0.5}$ in most cases, significantly improving degraded density maps and enhancing the resolution of helices and loops, as illustrated in Fig. 3. Notably, the FSC curves demonstrate that CryoFM successfully improves the signal across all frequencies, while the baselines fails to retain trustworthy signals in the low-frequency range.
@@ 5.2 Anisotropic Noise Denoising
Figure 4: Forward operators for (i) the anisotropic noise and (ii) the missing wedge.
Anisotropic noise in cryo-EM occurs when noise distribution varies by direction, affecting some orientations more than others and leading to uneven reconstruction quality. As a result, certain orientations have lower signal-to-noise ratios than others. To approximate this degradation in a simplified manner, we amplify the spectral noise by a factor when the particle orientation falls within a specific range of angles, simulating the increased uncertainty in less-sampled orientations. The forward model $\mathcal{A}:\mathbb{C}^{D\times D\times D}\to\mathbb{C}^{D\times D\times D}$ is given by:
@@
where ${\mathbf{\epsilon}}\in\mathbb{R}^{D\times D\times D}$ is the same as defined in Eq. (6), and $\alpha>1$ is a scalar that increases the noise for specific orientation angles, reflecting the heightened uncertainty in those directions. $\theta_{\min}$ and $\theta_{\max}$ are the angles that controls the portion of the noise being amplified as illustrated in Fig. 4. In the experiment, we estimate $\sigma_{\text{noise}}^{2}(\nu)$ and $\alpha$ from the half maps. See Section F.2 for details.
@@
Figure 5: Result of the anisotropic noise denoising task. Two density maps from EMDB are added anisotropic spectral noise that the estimated resolution is 4.38 Å. The degraded density maps are filtered by relion_postprocess for visual clarity.
@@ 5.3 Missing Wedge Restoration
The missing wedge problem in cryo-ET arises from the limited tilt range of the electron microscope during data acquisition, causing incomplete Fourier space sampling and artifacts in the subtomograms. We simulate this effect by applying a wedge-shaped mask in the Fourier domain, removing data from unmeasured orientations. The forward model $\mathcal{A}:\mathbb{C}^{D\times D\times D}\to\mathbb{C}^{D\times D\times D}$ is given by:
@@
where $\theta_{\min}$ and $\theta_{\max}$ are typically set to $-60^{\circ}$ and $+60^{\circ}$ respectively, representing the common tilt angle limits in experimental setups as illustrated in Fig. 4. For posterior sampling, we slightly modify the flow posterior sampling algorithm to get Alg. 2. The modification is based on the fact that it is not necessary to compute the gradient of the loss function, since we can directly combine the observed part in ${\mathbf{y}}$ with the remaining part in ${\mathbf{x}}_{t}$ to maximize the likelihood.
@@
Figure 6: Result of the missing wedge restoration task. Two maps from EMDB are masked in Fourier space to simulate the missing wedge effect in cryo-ET. The central slices through the maps are also shown.
@@ 5.4 Ab initio Modeling
Ab initio modeling in cryo-EM involves reconstructing a coarse density map from 2D particle projections, which serves as a reference for refinement. Here, we simplify the problem by focusing on generating a coarse density map from a few clean 2D projections obtained via upstream 2D classification. This approach mirrors earlier ab initio methods (Ludtke et al., 1999; Voss et al., 2010) before cryoSPARC’s introduction of SGD (Punjani et al., 2017). Given $K$ projections, there exist $K$ forward operators, where the $k$-th forward model is $\mathcal{A}^{(k)}:\mathbb{R}^{D\times D\times D}\to\mathbb{R}^{D\times D}$ given by:
@@
where $\mathcal{P}$ is a projection operator in the real space, and $\phi^{(k)}\in\mathbb{SO}(3)\times\mathbb{R}^{2}$ is the pose of the 2D projections. In this task, $\phi^{(k)}$ is unknown and can not be pre-determined. We modify the flow posterior sampling to get Alg. 3, where the main modifications are: (i) searching the optimal pose iteratively in the sampling process (Scheres, 2012a; Punjani et al., 2017; Zhong et al., 2021b), and (ii) computing the likelihood using correlation to avoid numerical issues. Moreover, since ab initio modeling focuses on capturing the global shape at low resolution, we use CryoFM-L for this task.